I need help filling out this comparative advantage chart.Opportunity Cost Agriculture Industrial Industrial Units produced by one worker in one year: Workforce (people) Agriculture 1
Answer #1

1) Opportunity cost of producing agriculture in China = units of industrial production/units of agriculture production = 10/10 = 1.

and opportunity cost of industrial production in China = units of agriculture production/units of industrial production = 10/10 = 1.

Similarly, opportunity cost of producing agriculture in USA = units of industrial production per worker/units of agriculture production per worker = 100/20 = 5.

and opportunity cost of industrial production in USA = units of agriculture production per worker/units of industrial production per worker = 20/100 = 0.2.

2) Initially, without trade, China's production = 1000*10 = 10,000 units of agriculture or 1000*10 = 10,000 units of industrial. China's PPF curve has vertical (Industrial) and horizontal (Agriculture) intercepts of 10,000 units each.

Let us assume, that China initially produces and consumes on its PPF curve, i.e, 5000 units of industrial and 5000 units of agriculture.

Similarly, without trade, USA's production = 300*20 = 6,000 units of agriculture or 300*100 = 30,000 units of industrial. Thus, USA's PPF curve has vertical (Industrial) and horizontal (Agriculture) intercepts of 30,000 units and 6,000 units respectively.

Let us assume, that USA initially produces and consumes on its PPF curve, i.e, 15,000 units of industrial and 3,000 units of agriculture.

With trade, China will only produce 10,000 units of agriculture and USA will only produce 30,000 units of industrial according to their individual comparative advantage.

A) With exchange rate of 3000 agriculture for 14000 industrial, China will consume 7000 units of agriculture and trades remaining 3000 units of agriculture to USA. In return it gains 14000 units of industrial. Thus, comparing to initial consumption of 5000 units each, China will gain 2000 units of agriculture and 9000 units of industrial from trade.

Similarly, with exchange rate of 3000 agriculture for 14000 industrial, USA will consume 16000 units of industrial and trades remaining 14000 units of industrial to China. In return it gains 3000 units of agriculture. Thus, comparing to initial consumption of 3000 units of agriculture and 15000 units of industrial, USA will gain 0 units of agriculture and 1000 units of industrial from trade.

B) Now, with exchange rate of 4000 agriculture for 6000 industrial, China will consume 6000 units of agriculture and trades remaining 4000 units of agriculture to USA. In return it gains 6000 units of industrial. Thus, comparing to initial consumption of 5000 units each, China will gain 1000 units of agriculture and 1000 units of industrial from trade.

Similarly, with exchange rate of 4000 agriculture for 6000 industrial, USA will consume 24000 units of industrial and trades remaining 6000 units of industrial to China. In return it gains 4000 units of agriculture. Thus, comparing to initial consumption of 3000 units of agriculture and 15000 units of industrial, USA will gain 1000 units of agriculture and 9000 units of industrial from trade.

The outcome without trade The outcome with trade ytrade What they consume what they produce what the produce we and consume (